OpenStudy (anonymous):

Solve the equation by completing the square , t^2-10t-24=0

OpenStudy (tkhunny):

Well, do it. What's first?

OpenStudy (anonymous):

The question is wrong it will be +24 :/

OpenStudy (tkhunny):

Not sure why you would know that, but it makes no difference for the methodology. Let's see it!

OpenStudy (anonymous):

that is what it says so

OpenStudy (anonymous):

i have four others like tht so im lost

OpenStudy (tkhunny):

It remains of no consequence. Please demonstrate the method. What's first? \(t^{2}-10t-24=0\) Needs to look like this: \((t-Something)^{2} = Something\;Else\)

OpenStudy (anonymous):

u have to set all up on the left side it says thats it i have no clue it for squares

OpenStudy (tkhunny):

You need only this principle \((a+b)^{2} = a^{2} + 2ab + b^{2}\)

OpenStudy (anonymous):

how u do this then

OpenStudy (dumbcow):

expanding on that principle ...say a is t \[(t+b)^{2} = t^{2} +2bt +b^{2}\] now ignore the constant (-24) for now \[t^{2} -10t + ? = t^{2} +2bt +b^{2}\] what does "b" need to be to make it fit (turn it into a perfect square) -10 = 2b --> b = -5 ---> b^2 = 25 now you have \[t^{2} -10t +25 = (t-5)^{2}\] thats completing the square ok to keep everything balanced, we need to subtract 25 since we added it to make the perfect square \[\rightarrow t^{2}-10t = (t^{2} -10t +25)-25 = (t-5)^{2} -25\] now add on the constant we ignored earlier \[\rightarrow (t-5)^{2} -25 -24 = 0\] now solve for t \[(t-5)^{2} = 49\]

OpenStudy (tkhunny):

Take a good, hard look at this: \((t-a)^{2} = t^{2} - 2at + a^2\) Now, take a good hard look at this: \(t^{2}-10t-24=0\)